# AI-Driven Predictive Analytics Approach for Early Prognosis of Chronic Kidney Disease Using Ensemble Learning and Explainable AI thanks:

### v1

@@ Abstract
Chronic Kidney Disease (CKD) is one of the widespread Chronic diseases with no known ultimo cure and high morbidity. Research demonstrates that progressive Chronic Kidney Disease (CKD) is a heterogeneous disorder that significantly impacts kidney structure and functions, eventually leading to kidney failure. With the progression of time, chronic kidney disease has moved from a life-threatening disease affecting few people to a common disorder of varying severity. The goal of this research is to visualize dominating features, feature scores, and values exhibited for early prognosis and detection of CKD using ensemble learning and explainable AI. For that, an AI-driven predictive analytics approach is proposed to aid clinical practitioners in prescribing lifestyle modifications for individual patients to reduce the rate of progression of this disease. Our dataset is collected on body vitals from individuals with CKD and healthy subjects to develop our proposed AI-driven solution accurately. In this regard, blood and urine test results are provided, and ensemble tree-based machine-learning models are applied to predict unseen cases of CKD. Our research findings are validated after lengthy consultations with nephrologists. Our experiments and interpretation results are compared with existing explainable AI applications in various healthcare domains, including CKD. The comparison shows that our developed AI models, particularly the Random Forest model, have identified more features as significant contributors than XgBoost. Interpretability (I), which measures the ratio of important to masked features, indicates that our XgBoost model achieved a higher score, specifically a Fidelity of 98%, in this metric and naturally in the FII index compared to competing models.
@@
Chronic Kidney Disease, Machine Learning Ensemble Models, Explainable AI, Interpretability
@@ I Introduction
The clinical definition of Chronic Kidney Disease is based on the presence of kidney damage or Decreased Kidney Function. That is when Glomerular Filtration Rate (GFR) less than 60 mL/min per 1.73 m2 for a period of 3 months or more. The risk factors of occurrence for CKD increase manifolds in the patients suffering from Diabetes Mellitus or Hypertension. Causation of the disease and management are measured in terms of stages of disease severity and clinical diagnosis. Routine Laboratory tests specifically Blood and Urine tests reveal the occurrence of CKD. The progression of the disease is also measured by GFR quantity. Management of this disease needs change in dietary plans and lifestyle modifications. But for all this to happen, it is highly important that the disease is detected beforehand. There are some factors that leads to kidney diseases like presence of Type I or Type II Diabetes, Presence of polycistic or other inherited kidney diseases, prolonged obstruction of urinary tract, recurrent kidney infection etc. The first sign of kidney disease from Diabetes is protein in the urine. When the filters are damaged, usually a protein named Albumin passes into the urine from the blood. But a healthy kidney won’t let this occur. High blood pressure can cause CKD and can happen as a result of CKD as well [1]. According to reports of Centers of Disease Control and Prevention (CDC), about 37 Million adults in USA are unaware about having CKD. In the US, approximately 360 people start dialysis treatment for kidney failure every hour. Almost three out of four new cases of kidney failure occurs due to high blood pressure and diabetes which makes them leading causes of kidney failure in the United States [2].
Chronic Kidney Disease is a severe public health problem that affects general population of all ages. According to reports of Mayo Clinic, Chronic Kidney Disease (CKD) could also lead to Anemia, Fluid Retention, High Blood Pressure, Weak Bones, Risk of Bone Fractures, Pregnancy Complications and cause decreased Immunoresponse [3]. With Chronic Kidney Disease, the kidneys of patients cannot make enough Erythropoietin which causes drop in red blood cells and gradually leads to Anemia. Kidneys are the primary organ for filtering waste materials. When that organ loses it’s capacity of filtering, naturally fluid overload, electrolytes and wastes are built up on the body. Furthermore, CKD causes the blood vessels to get narrowed which increases Blood Pressure. CKD causes vitamin D synthesis leading to Weak Bones[3]. More symptoms of CKD are Nausea, Vomiting, Loss of Appetite, Muscle cramps, Swelling of Feet and Ankles, Shortness of Breath etc. It is hard to identify between the causes and complications of CKD without clinical diagnosis and follow ups. Because complication of CKD can also be a cause and vice versa.
@@
Chronic Kidney Disease (CKD) is a major health concern that affects millions of people around the world every year. The irreversible nature of this disease not just leads to comorbid diseases like Diabetes Mellitus, Hypertension etc. but also it can permanently damage the kidney by progressing to Acute Kidney Injury (AKI) or End Stage Renal Diseases (ESRD). Unlike the healthcare facilities available at developed countries, in the developing countries, the CKD is not identified till the ESRD occurs where dialysis remains the only option. The common practice associated with identification of CKD is if the Glomerular Filtration Rate (GFR) quantity is less than 60 mL/min or presence of biomarkers that indicate kidney damage for last 3 months. In most of the clinical practices, common tests for CKD includes measurement of GFR. GFR quantity is calculated from the serum creatinine concentration. From the Urine test, albuminuria or albumin to creatinine ratio is indicative of CKD [3].
@@
Tackling chronic kidney diseases has many economical challenges also. According to reports from American Society of Nephrology, approximately 26 million Americans have some evidence of chronic kidney disease and are at risk to develop kidney failure. Another 20 million are at risk of developing the kidney diseases. The number of people diagnosed with CKD has doubled in the past two decades and older Americans in the age range of 60-70 are more at risk of developing this disease. The progression of kidney disease eventually leads to End Stage Renal Diseases (ESRD) and the annual cost of treating ESRD is currently over \$32 billion. To put things more into global perspective, chronic kidney disease ranks 12th as a cause of death and 17th as a cause of disability [4]. With the state of the art healthcare facilities in the West, less reports are obtained for developing countries in the South-East Asia. In the public hospitals of the Indian sub-continent, treatment of ESRD has become a low priority. With the absence of health insurance plans, less than 10% of ESRD patients receive any Renal Replacement Therapy. Treatment gets stopped for vast majority of the patients in first 3 months due to the overwhelming cost of the process[4].
@@
Among existing literature on the applications of XAI on chronic kidney disease, studies conducted by Manju et al at [66] have used traditional pre-processing techniques like min max scaling and Z-score standardization to keep all the features in a range. Z-score standardization usually would transform the values of features so that the standard deviation of features would be 1 with the mean value 0. If x is the actual value of the feature and x’ is the transformed value of the feature after Z-score standardization, then
@@ II Background Study
II-A Understanding Causes and Effects of CKD
The advanced stage of chronic kidney disease (CKD), known as end-stage renal disease (ESRD), signifies a critical point where patients require either dialysis or a kidney transplant. The stakes of prevention of CKD towards ESRD become even higher considering the numbers game. In the United States, approximately 20 Million people are suffering from CKD. This number would only increase with the aging of the States population. The economic aspect of tackling CKD is further dangerous as nearly USD 28 Billion is spent every year on ESRD in the United States alone. The rising prevalence of different kinds of kidney diseases is majorly concerning for healthcare professionals and patients. Nearly 8 Million people which is roughly 4% of the entire population of the US have moderate to severe form of CKD and another 450,000 people have ESRD [65].
Identifying, diagnosing, and managing the diverse causes of CKD often experience delays. Since the causes and results of CKD are kind of similar, it is hard to distinguish from the perspective of a layman why CKD occurs and what they need to follow in order to prevent it. This is where AI and Machine Learning models come to good use to boil down the cause and effects of CKD. Analyzing the literature review, most of the existing work on CKD with Machine Learning focused on areas like CKD prognosis, CKD diagnosis, Risks of CKD development, CKD treatment etc. And these research studies’ commonly used variables are Age and Gender (Demographic), Blood Pressure, Haemoglobin, Serum Creatinine (Clinical) [5].
II-B Dealing with Healthcare Datasets
Getting the hands on healthcare datasets in itself is a difficult task. The datasets themselves are very rarely optimized for Machine Learning applications. Most of the clinical studies are longitudinal cohort studies where the subjects are regularly followed up after certain intervals. It is natural in these datasets to have missing values in certain features as many subjects from these studies chose not to follow up anymore. Moreover, some cross-sectional clinical data are prone to human error on the data entry. The missing data occurrence in these studies add more complexity to the already complicated domain of healthcare. Several studies conducted by recently have used iterative imputation techniques to account for missing value imputation, and used naive supervised learning techniques to predict CKD cases validated by cross validations on the test dataset [5]. Among other studies specially on the CKD diagnosis from patient data from Brazil, they addressed some critical aspects specially for early detection of CKD [6]. In the developing countries like Brazil, it becomes very important to aid the physicians from an automated system because in the clinical settings, Kidney Diseases are not detected specially till they reach towards ESRD. The requests for the clinical datasets have to go through proper channels, authorities. The researchers requiring these datasets also have to follow the code of compliance dealing with human subjects and perform training procedures before diving into research with the clinical data.
@@ II-C Application of XAI in CKD Research
Among other XAI applications in different domains of healthcare, studies conducted by Devam Dave et al use example based techniques along with feature based techniques [11]. They used Anchors which generate local regions that denotes a better construction of the dataset for explanation and also have less computational expense than SHAP. Anchors were used to identify as “if then” conditionals with independent feature values as if condition and the dependent feature value as “then” statement. The Counterfactuals were employed to evaluate what sort of changes would need to occur in the independent features so that the prediction of the model is switched to the contrasting class. Counterfactuals with prototypes provide similar but faster explanations on the basis of prototype or samples that is representative of the instances belonging to a class. In the contrastive explanation models, Pertinent Positives work similar to Anchors where Pertinent Negatives work similar to Counterfactual [12]. The concept of integrated gradients from [13] has been explained pretty well from the MNIST Fashion dataset at [11]. For the Integrated gradients, positive attributions are the ones that make a positive influence for the model to identify it as the positive class and negative attributions are the opposite ones. They provided the example with “Shoes” being the positive class where pixel values of features like laces and shoes were the positive attributions and some features of shoe laces contributed towards identifying it as “Not a Shoe” or the negative class.
II-D Addressing Data Security in AI Applications
@@
TABLE I: Percentage of Missingness in all columns
@@ III-A Missing Values
Since the dataset is collected from manual input of the patients vitals and results of blood tests, naturally there are lots of missing values in the dataset. From table I, it is pretty evident that 12 out of the 25 total features had moderate to high amount of missing values.
@@
TABLE II: RJA Little’s MCAR Test On Different Sample Sizes
@@
Experiments were designed to analyze whether this null hypothesis could be rejected or not. This research used the extensive test conducted by RJA Little at [63] where the significance level, degrees of freedom and test statistic P-value would be analyzed to find whether the null hypothesis could be rejected or not. The significance level was determined by Fisher’s threshold [64]. The Little’s MCAR Test needs all the variables to be numerical so the categorical variables were converted to numeric ones and then randomly 20%, 30% and 40% samples were chosen for the test.
@@
TABLE III: Performance Evaluation of
the Imputation Techniques
Technique Evaluation Algorithm RMSE
Mean Imputation Logistic Regression 0.46
Median Imputation Logistic Regression 0.45
KNN (K = 2) Logistic Regression 0.45
KNN (K = 3) Logistic Regression 0.55
KNN (K = 5) Logistic Regression 0.50
Mean Imputation Naive Bayes 0.22
Median Imputation Naive Bayes 0.31
KNN (K = 3) Naive Bayes 0.31
KNN (K = 3) Naive Bayes 0.22
KNN (K = 3) Naive Bayes 0.23
TABLE IV: Highly Correlating Features with Target Feature
@@ III-B Imputation of Missing Values
Caution was adapted in dealing with missing values in this dataset since it is a sensitive domain of patient health records. Firstly traditional Mean Median modes of missing value imputation was adapted. It was necessary to validate our decision making in every step so that all the processes followed a well defined and calculative approach. So, the performance of the Machine Learning models in the dataset with missing values was evaluated, then the same models on the imputed dataset was fitted with the mean and median values. In order to understand the quality of imputation of the missing values in such a sensitive domain, the Root Mean Square Error (RMSE) error metric was employed for all the imputation techniques. The imputation results with the ML models are described in the following table:
From the table III, it is evident that Mean Imputation and K-Nearest Neighbour methods of missing value imputations have yielded the best results with the lowest RMSE values on this dataset. For K-Nearest Neighbor (KNN) imputation, trial and error method was employed where values greater than 5 had significantly higher RMSE values hence those are not reported.

### v2

@@ Abstract
Chronic Kidney Disease (CKD) is one of the widespread Chronic diseases with no known ultimo cure and high morbidity. Research demonstrates that progressive CKD is a heterogeneous disorder that significantly impacts kidney structure and functions, eventually leading to kidney failure. The goal of this research is to first develop an accurate ensemble model for prediction of unseen cases of CKD given the biomarkers. Also, we have implemented the Explainable AI (XAI) algorithms to interpret the decision-making process of the ensemble models in terms of dominating features and the feature values. The takeaway from our research is to aid the physicians make an informed decision about the disease and provide a case by case explanation behind their decisions. Also, XAI algorithms would allow the patients or subjects understand the causes behind their disease at early stages so that they can be cautious about the progression of the disease to later stages.
@@
CKD, Machine Learning Ensemble Models, Explainable AI, Interpretability
[…]
=-21pt
@@ I Introduction
The clinical definition of CKD is based on the presence of kidney damage or Decreased Kidney Function. That is when Glomerular Filtration Rate (GFR) less than 60 mL/min per 1.73 m2 for a period of 3 months or more.
I-A Analysis of biomarkers for CKD
@@
The risk factors of occurrence for CKD increase manifolds in the patients suffering from Diabetes Mellitus or Hypertension. Causation of the disease and management are measured in terms of stages of disease severity and clinical diagnosis. Routine Laboratory tests specifically Blood and Urine tests reveal the occurrence of CKD. The progression of the disease is also measured by GFR quantity. Management of this disease needs change in dietary plans and lifestyle modifications. But for all this to happen, it is highly important that the disease is detected beforehand. There are some factors that leads to kidney diseases like presence of Type I or Type II Diabetes, Presence of polycistic or other inherited kidney diseases, prolonged obstruction of urinary tract, recurrent kidney infection etc. The first sign of kidney disease from Diabetes is protein in the urine. When the filters are damaged, usually a protein named Albumin passes into the urine from the blood. But a healthy kidney will not let this occur. High blood pressure can cause CKD and can happen as a result of CKD as well [1]. CKD is a major health concern that affects millions of people around the world every year. The irreversible nature of this disease not just leads to comorbid diseases like Diabetes Mellitus, Hypertension etc. but also it can permanently damage the kidney by progressing to Acute Kidney Injury (AKI) or End Stage Renal Diseases (ESRD). Unlike the healthcare facilities available at developed countries, in the developing countries, the CKD is not identified till the ESRD occurs where dialysis remains the only option. The common practice associated with identification of CKD is if the Glomerular Filtration Rate (GFR) quantity is less than 60 mL/min or presence of biomarkers that indicate kidney damage for last 3 months. In most of the clinical practices, common tests for CKD includes measurement of GFR. GFR quantity is calculated from the serum creatinine concentration. From the Urine test, albuminuria or albumin to creatinine ratio is indicative of CKD [3].
I-B CKD as a Health Concern
CKD is a severe public health problem that affects general population of all ages. According to reports of Centers of Disease Control and Prevention (CDC), about 37 Million adults in USA are unaware about having CKD. In the US, approximately 360 people start dialysis treatment for kidney failure every hour. Almost three out of four new cases of kidney failure occurs due to high blood pressure and diabetes which makes them leading causes of kidney failure in the United States [2]. According to reports of Mayo Clinic, CKD could also lead to Anemia, Fluid Retention, High Blood Pressure, Weak Bones, Risk of Bone Fractures, Pregnancy Complications and cause decreased Immunoresponse [3]. With CKD, the kidneys of patients cannot make enough Erythropoietin which causes drop in red blood cells and gradually leads to Anemia. Kidneys are the primary organ for filtering waste materials. When that organ loses it’s capacity of filtering, naturally fluid overload, electrolytes and wastes are built up on the body. Furthermore, CKD causes the blood vessels to get narrowed which increases Blood Pressure. CKD causes vitamin D synthesis leading to Weak Bones[3]. More symptoms of CKD are Nausea, Vomiting, Loss of Appetite, Muscle cramps, Swelling of Feet and Ankles, Shortness of Breath etc. It is hard to identify between the causes and complications of CKD without clinical diagnosis and follow ups. Because complication of CKD can also be a cause and vice versa.
I-C CKD in Monetary Numbers
@@
Tackling chronic kidney diseases has many economical challenges also. According to reports from American Society of Nephrology, approximately 26 million Americans have some evidence of CKD and are at risk to develop kidney failure. Another 20 million are at risk of developing the kidney diseases. The number of people diagnosed with CKD has doubled in the past two decades and older Americans in the age range of 60-70 are more at risk of developing this disease. The progression of kidney disease eventually leads to End Stage Renal Diseases (ESRD) and the annual cost of treating ESRD is currently over \$32 billion. To put things more into global perspective, CKD ranks 12th as a cause of death and 17th as a cause of disability [4]. With the state of the art healthcare facilities in the West, less reports are obtained for developing countries in the South-East Asia. In the public hospitals of the Indian sub-continent, treatment of ESRD has become a low priority. With the absence of health insurance plans, less than 10% of ESRD patients receive any Renal Replacement Therapy. Treatment gets stopped for vast majority of the patients in first 3 months due to the overwhelming cost of the process[4].
I-D Difficulty in CKD Analysis from Engineering Perspective
@@ I-E A Brief on Existing Research on CKD
I-F Use of XAI in CKD
Among existing literature on the applications of XAI on CKD, studies conducted by Manju et al at [66] have used traditional pre-processing techniques like min max scaling and Z-score standardization to keep all the features in a range. Z-score standardization usually would transform the values of features so that the standard deviation of features would be 1 with the mean value 0. If x is the actual value of the feature and x’ is the transformed value of the feature after Z-score standardization, then
@@ II Background Study
II-A Dealing with Healthcare Datasets
Most of the clinical studies are longitudinal cohort studies where the subjects are regularly followed up after certain intervals. It is natural in these datasets to have missing values in certain features as many subjects from these studies chose not to follow up anymore. Moreover, some cross-sectional clinical data are prone to human error on the data entry. The missing data occurrence in these studies add more complexity to the already complicated domain of healthcare. Several studies conducted by recently have used iterative imputation techniques to account for missing value imputation, and used naive supervised learning techniques to predict CKD cases validated by cross validations on the test dataset [5]. Among other studies specially on the CKD diagnosis from patient data from Brazil, they addressed some critical aspects specially for early detection of CKD [6]. In the developing countries like Brazil, it becomes very important to aid the physicians from an automated system because in the clinical settings, Kidney Diseases are not detected specially till they reach towards ESRD. The requests for the clinical datasets have to go through proper channels, authorities. The researchers requiring these datasets also have to follow the code of compliance dealing with human subjects and perform training procedures before diving into research with the clinical data.
@@ II-B Application of XAI in CKD Research
Among other XAI applications in different domains of healthcare, studies conducted by Devam Dave et al use example based techniques along with feature based techniques [11]. They used Anchors which generate local regions that denotes a better construction of the dataset for explanation and also have less computational expense than SHAP. Anchors were used to identify as “if then” conditionals with independent feature values as if condition and the dependent feature value as “then” statement. The Counterfactuals were employed to evaluate what sort of changes would need to occur in the independent features so that the prediction of the model is switched to the contrasting class. Counterfactuals with prototypes provide similar but faster explanations on the basis of prototype or samples that is representative of the instances belonging to a class. In the contrastive explanation models, Pertinent Positives work similar to Anchors where Pertinent Negatives work similar to Counterfactual [12]. The concept of integrated gradients from [13] has been explained pretty well from the MNIST Fashion dataset at [11]. For the Integrated gradients, positive attributions are the ones that make a positive influence for the model to identify it as the positive class and negative attributions are the opposite ones. They provided the example with “Shoes” being the positive class where pixel values of features like laces and shoes were the positive attributions and some features of shoe laces contributed towards identifying it as “Not a Shoe” or the negative class. Tables 1 contains different techniques employed by authors of the background study with their pros and cons. Table 2 contains a more detailed analysis of the specific dataset used, experiment strategies, models used and the accuracy metric results.
TABLE I: Comparison of Techniques for CKD Prediction
Authors Techniques Pros Cons
Silveira et al [6] Data augmentation by consultation with nephrologist, SMOTE. Both techniques are used for handling class imbalanced dataset. Use of Grid Search Cross Validation to find best set of parameters for the ensemble models increasing time complexity.
Vintrella et al [7] Collection of patient’s demographic information. Family history information gathering like family history of diabetes, anemia, hypertension, presence or absence of kidney stones, smoking etc. Research analysis limited to only Caucasian ethnicity. They have not used features like Body Mass Index (BMI), Protein Urea, Diet etc. which are important contributors in CKD prediction.
Pedro et al [8] XAI feature based techniques on ensemble models. Calculation of interpretive measures like Interpretability (I), Fidelity (F). Lack of validation data as research is limited to analysis of one CKD dataset from UCI ML repository.
Bayram et al [9] Use of XAI on Kidney Image Dataset. Distinction between different kidney diseases like kidney cyst, kidney stone by XAI. Incorrect predictions were not analyzed, dataset was limited to only one ethnicity.
Gabriel et al [10] Case Based Reasoning (CBR) on 20,000 patients with 7,493 variables encompassing demographic information and previous diagnosis of CKD. Neural Net Based CBR was employed to process the explanation of the CKD decision based on similar examples seen earlier. Selected features in their experiments were binary, they have not worked on any continuous variable. User feedback from the explanatory models is not implemented.
Devam et al [19] Example Based XAI techniques on Heart Disease Dataset. Use of Anchors as “If-then” statement and Counter-factuals to understand the changes need to happen for each feature to predict the contrasting class. Comparative Analysis not found with concurrent datasets on Heart Disease or other medical datasets.
Our methods Missing data analysis by JA Little’s MCAR Test, Imputation by ML models and nephrologists, Feature and Example based XAI applications. Thorough analysis of missing data, precise feature selection methods, patient specific contrastive class prediction details, rigorous pre-processing of dataset, application of novel fidelity measures on this CKD dataset, approaches of building a more clinically comprehensive CKD dataset. No discussions on demographic variables, no calculation of internal fidelity, less analysis of inter-feature correlation in independent features.
TABLE II: Experimental strategies and models in CKD prediction studies.
Study Title Dataset Used Experiment Strategy Models Used Accuracy Precision Recall F1 Score
Silveira et al [6] Medical records of Brazilians with or without CKD, having attributes like hypertension, diabetes, creatinine, etc. Hold-out validation, multiple stratified cross-validation (CV), and nested cross-validation. Decision Tree (DT), Random Forest, and Multi-Class AdaBoosted DTs 0.95 0.92 0.93 0.94
Vintrella et al [7] Data extracted from the Vimercate Hospital EMR with features like GFR over a time period, diabetes, chlorine, cardiopathic, anemic, etc. Classification and regression approaches to determine exactly when patients would need dialysis. For classification metrics, 10-fold cross-validation was used to analyze accuracy metrics. For regression, Mean Squared Error (MSE) was used to estimate prediction error. Logistic Regression, Decision Tree (DT), Random Forest, Fully Connected Neural Nets 0.95 0.92 0.93 0.94
Pedro et al [8] CKD dataset from the UCI ML Repository. 5-fold cross-validation used to measure unseen test set predictions, with different interpretability measures calculated using permutation feature importance techniques. Random Forest, Decision Tree (DT), AdaBoost, Extra Trees Classifier 0.95 0.92 0.93 0.94
Bayram et al [9] The kidney dataset containing images of patients diagnosed with kidney cysts, normal, and stone findings. A supervised image classification task using deep learning models. You Only Look Once (YOLO v7) and YOLO v7 Tiny architectures 0.85 0.88 0.83 0.85
Gabriel et al [10] RIPS database collected from the Colombian Ministry of Health and Social Protection that contains individual record of health service delivery. Neural network with hyperparameters tuned on the dataset. To tackle overfitting, early stopping and dropout were applied as regularization techniques. Neural Network, Support Vector Machine (SVM), and Random Forest (RF) 0.95 0.94 0.97 0.95
Devam et al [19] The heart disease dataset from Cleveland UC Irvine dataset. Feature-based and example-based XAI algorithms are applied on the ensemble models for prediction analysis. Logistic Regression (LR), XGBoost; SHAP, Counterfactuals, Anchors, Integrated Gradients applied on LR and XGBoost N/A N/A N/A N/A
II-C Addressing Data Security in AI Applications
@@
TABLE III: Percentage of Missingness in all columns
@@ III-A Missing Values
Since the dataset is collected from manual input of the patients vitals and results of blood tests, naturally there are lots of missing values in the dataset. From table I, it is observed that 12 out of the 25 total features had moderate to high amount of missing values.
@@
TABLE IV: RJA Little’s MCAR Test On Different Sample Sizes
@@
Experiments were designed to analyze whether this null hypothesis could be rejected or not. This research used the extensive test conducted by RJA Little at [63] where the significance level, degrees of freedom and test statistic P-value would be analyzed to find whether the null hypothesis could be rejected or not. The significance level was determined by Fisher’s threshold [64]. The Little’s MCAR Test needs all the variables to be numerical so the categorical variables were converted to numeric ones and then randomly 30%, 60% and 90% samples were chosen for the test.
@@
TABLE V: Highly Correlating Features with Target Feature
@@ III-B Imputation of Missing Values
The dataset had a significant percentage of missing values in a number of columns. Previous approaches of mean, median or KNN imputations were discarded because it could lead to data leakage issue of the subjects. Moreover, in a sensitive issue like medical datasets, just using traditional means of imputation is not quite clinically feasible. So, a more effective measure was using training-testing split of the original dataset to impute the missing values. First we found out the exact rows where none of the features had any missing values and used that particular dataset as the training dataset. Training was performed with a linear regression model as we had to predict the continuous values in some features. Then we did some pre-processing like cleaning strings, tabs, whitespaces on the test dataset with the missing values. One key issue in the missing data imputation was it was difficult to select a target feature and treat other features as independent in the test dataset. This happened because whichever feature we chose as dependent, there were more than a couple of independent features which had missing values. This is where we had to sort the features in the test dataset by choosing one feature as target feature for imputation. If that target feature was termed as ”A”, we looked at which are the exact features that had values present corresponding to the missing values of feature ”A”. Then in the training dataset, we just trained on those set of features to impute the values in feature ”A”. Similar process was followed for imputation of all the features that had missing values. Some post-processing was also required to organize the imputed values in the final imputed dataset. The data imputation process was also validated by the nephrologists Dr. Md. Nabiul Hassan Rana from Asgar Ali Medical College, Dhaka, Bangladesh and Dr. Nazneen Mahmood from Anwar Khan Modern Hospital, Dhaka. Both of them agreed that the mathematical mean, median imputation on clinical data could lead to wrong calculations and introduce bias in the models. So, after a collective discussion, they also agreed that the best way to impute the missing data in this clinical setting would be to look at the observed values, learn the representations and then use the learnt representations to get an estimate of the missing data.